what is my speed if I travel 18 miles in 50 minutes
Answer by oberobic(2304)

d = r*t
d = 18 miles
.
t = 50 min
t = 5/6 hr
.
r = d/t
r = 18/(5/6)
r = 21.6 miles/hr
.
Done.